A simple random sample of 100 postal employees is used to test if the average time postal employees have worked for the postal service has changed from the value of 7.5 years recorded 20 years ago. The sample mean was 7 years with a standard deviation of s = 2 years. Assume the distribution of the time the employees have worked for the postal services is approximately normal.

What are the appropriate degrees of freedom (df) for this test?

100

7

19

48

99
Answer #1

solution

given that

n =100

Degrees of freedom = df = n - 1 = 100- 1 = 99

correct answer (df)=99